# Modelling Global Trade with Optimal Transport

### v1

@@ Results
   [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] : Figure 7Comparison with gravity model. A–B Comparison of the RMSE accuracies of the estimated transport volumes of the OT approach (blue) and the gravity model (orange). Values are averaged over all countries and years, with the errorbars showing one standard deviation from the mean (triangular marker). Also shown are the median values (diamond markers). Shown is the RMSE (left) and the RMSE in units of the standard deviation on the true data (right). C–D Comparison plot of the OT and gravity estimates ($x$-axis) versus the true data ($y$-axis), for the two commodities on which the gravity model performs best. Also shown is a linear fit (dotted line), its estimated slope $m$, the Pearson coefficient of the fit $r$, and the line $y=x$ (solid line). See fig. LABEL:fig:Gravity_comparisons_all in the appendix for an overview of all commodities.
@@
Lastly, we compare the performance of our method with a standard gravity model [11, 36, 37, 38], as specified in equations (1)–(2). The covariates include geographic distance, shared borders, colonial ties, common language, regional trade agreements, tariffs, and importer/exporter fixed effects to account for multilateral resistance, resulting in a $m$+$n$+11-dimensional estimation problem (see Supplementary Information for details). We estimate the coefficients using least-squares optimization and compare the accuracy of the estimated transport plans $\mathbf{T}$. Figure LABEL:fig:Gravity_compA shows that the OT approach significantly outperforms the gravity model in terms of RMSE, often by two to three orders of magnitude. Figure LABEL:fig:Gravity_compB illustrates that OT estimates typically fall within one standard deviation of the data uncertainty, whereas gravity estimates tend to range from one to two, at times even three to four standard deviations. The gravity model also exhibits much higher variance in accuracy compared to OT. Figures LABEL:fig:Gravity_compC–D present a scatter plot of OT (blue) and gravity (orange) estimates against FAO data. For the two commodities where the gravity model performs best, a linear fit through the OT estimates yields a near-perfect slope of $m=1$ with a Pearson coefficient close to 1, perfectly matching the tail end of the distribution. In contrast, the gravity model struggles, especially in the upper tail, likely due to model misspecification.
@@ Discussion
The classic OT problem eq. (6) can be interpreted as the central’s planners problem of finding the optimal assigment/matching of supplies and demands. The dual OT problem is given by
@@
All code and data is available at https://github.com/ThGaskin/NeuralABM. Instructions for running the model is given in the README.
@@ Acknowledgements
$+\lambda_{7}\mathrm{RTA}_{ij}+\lambda_{8}\log\pi_{i}+\lambda_{9}\log\chi_{j}+\lambda_{10}\log\mathrm{TRFF}_{ij}+\lambda_{11}\mathrm{NTM}_{ij}.$ (20)
@@
    $\kappa_{i}$ are the exporter-fixed effects
@@
    $\omega_{j}$ are the importer-fixed effects
@@
    $E_{j}$ is the total consumption of the importer, in tonnes
@@
    $\pi_{i}=\sum_{j}d_{ij}E_{j}/\sum_{k}E_{k}$ is the remoteness index of the exporter,
@@
  13. 13.
    NTM is the number of non-tariff trade restrictions, as given by the UNCTAD TRAINS dataset [37]. NTMij is simply the total number of trade restrictions in place.
The remoteness indices account for the multilateral resistances [11]. This gives a $m+n+11$-dimensional regression problem for each commodity. We estimate the parameters $\{\lambda_{k}\}$ using a least-squares optimisation. Table LABEL:tab:Gravity_parameters gives the estimated parameters for each commodity. Figure LABEL:fig:Gravity_comparisons_all plots the estimated values $\hat{T}_{ij}$ against the reporter-averaged FAOStat values for both the OT and the gravity models. Also shown are a linear fit with slopes and Pearson coefficients indicated.

### v2

@@ Results
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] : Figure 7Comparison with gravity model. A–B Comparison plot of the OT and gravity estimates ($x$-axis) versus the true data ($y$-axis) on two selected commodities. Also shown is a linear fit (dotted line), its estimated slope $m$, the Pearson coefficient of the fit $r$, and the line $y=x$ (solid line). See fig. LABEL:fig:Gravity_comparisons_all in the appendix for an overview of all commodities. C–D Comparison of the RMSE accuracies of the estimated transport volumes of the OT approach (blue) and the gravity model (orange). Values are averaged over all countries and years, with the errorbars showing one standard deviation from the mean (triangular marker). Also shown are the median values (diamond markers). Shown are the RMSE (left) and the RMSE in units of the standard deviation on the true data (right).
@@
Lastly, we compare the performance of our method with a standard gravity model [11, 36, 37, 38], as specified in equations (1)–(2). The covariates include geographic distance, shared borders, colonial ties, common language, regional trade agreements, tariffs, and importer/exporter fixed effects to account for multilateral resistance (see Supplementary Information for details). We estimate the coefficients using Poisson Pseudo Maximum Likelihood estimation and compare the accuracy of the estimated transport plans $\mathbf{T}$. Figures LABEL:fig:Gravity_compA–B show scatter plots of the OT (blue) and gravity (orange) estimates against the FAO data. For all commodities studied, a linear fit through the OT estimates yields a near-perfect slope of $m=1$ with a Pearson coefficient close to 1, perfectly fitting the tail end of the distribution. In contrast, the gravity model’s performance is much more volatile, with linear fits ranging from a Pearson coefficient of between 0.998 (best) to 0.747 (worst) (see also Supplementary Information). Due to model misspecification, the fits to the tails of the distributions are generally significantly poorer. Consequently, figure LABEL:fig:Gravity_compC shows that the OT approach significantly outperforms the gravity model in terms of RMSE, often by two to three orders of magnitude. Figure LABEL:fig:Gravity_compD illustrates that OT estimates typically fall within one standard deviation of the data uncertainty, whereas gravity estimates tend to range from one to two, at times even three to four standard deviations. The gravity model also exhibits much higher variance in accuracy compared to OT.
@@ Discussion
The classic OT problem eq. (6) can be interpreted as the central’s planners problem of finding the optimal assignment/matching of supplies and demands. The dual OT problem is given by
@@
All code and data is available at https://github.com/ThGaskin/inverse-optimal-transport. Instructions for running the model are given in the README.
@@ Acknowledgements
$+\lambda_{7}\mathrm{RTA}_{ij}+\lambda_{8}\log\chi_{j}+\lambda_{9}\log\mathrm{TRFF}_{ij}.$ (20)
@@
    $\kappa_{i}$ are the time-dependent exporter-fixed effects,
@@
    $\omega_{j}$ are the time-dependent importer-fixed effects,
@@
    $E_{j}$ is the total consumption of the importer, in tonnes,
@@
    $\chi_{j}=\sum_{i}d_{ij}O_{i}/\sum_{k}O_{k}$ is the remoteness index of the importer,
@@
The remoteness index accounts for the multilateral resistances [11]. This gives a $L\times(m+n)+9$-dimensional regression problem for each commodity, where $L=22$ are the number of years in the dataset (the fixed-effects are time-dependent). We estimate the parameters $\{\lambda_{k}\}$ using Poisson Pseudo Maximum Likelihood optimisation. Table LABEL:tab:Gravity_parameters gives the estimated parameters for each commodity. Figure LABEL:fig:Gravity_comparisons_all plots the estimated values $\hat{T}_{ij}$ against the reporter-averaged FAOStat values for both the OT and the gravity models. Also shown are a linear fit with slopes and Pearson coefficients indicated.